OpenStudy (anonymous):

{(-5, -15), (-3, -9), (0, 0) , (2, 6) , (4, 12)} Identify the rule for the function. The range is equal to four more than the domain. The range is equal to eight less than the domain. The range is equal to three more than the domain. The range is equal to three times the domain.

OpenStudy (anonymous):

Could i find someone ive asked questions on multiple accounts and not once an answer :/

jimthompson5910 (jim_thompson5910):

hint: the equation is of the form y = k*x

jimthompson5910 (jim_thompson5910):

where k is some fixed number

jimthompson5910 (jim_thompson5910):

Also remember that domain = set of possible inputs (ie x values) range = set of possible outputs (ie y values)

OpenStudy (anonymous):

im still lost

jimthompson5910 (jim_thompson5910):

pick any ordered pair you want that's not (0,0)

OpenStudy (anonymous):

-5,15

jimthompson5910 (jim_thompson5910):

so x = -5, y = 15

jimthompson5910 (jim_thompson5910):

plug those values into y = k*x then solve for k

jimthompson5910 (jim_thompson5910):

sry, y = -15

OpenStudy (jdoe0001):

\(\begin{array}{rrrllll} x&y\\ \hline\\ -5& -15\implies -5\times 3\\ -3& -9\implies -3\times 3\\ 0& 0\implies 0\times 3\\ 2& 6\implies 2\times 3\\ 4& 12\implies 4\times 3 \end{array}\)

OpenStudy (anonymous):

and x is -5 right?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

giving you like what -60?

jimthompson5910 (jim_thompson5910):

y = k*x y = k*(-5) ... plug in x = -5 -15 = k*(-5) ... plug in y = -15 now solve for k

OpenStudy (anonymous):

i feel really mental but im still lost .-.

jimthompson5910 (jim_thompson5910):

k*(-5) means "k times -5" so how do you undo multiplication?

OpenStudy (anonymous):

division? :p

jimthompson5910 (jim_thompson5910):

so you divide both sides by -5 to isolate k

OpenStudy (anonymous):

1 sec grabbing note book 2 work this out

OpenStudy (anonymous):

meh

jimthompson5910 (jim_thompson5910):

-15 divided by -5 = ???

OpenStudy (anonymous):

-3 :p

jimthompson5910 (jim_thompson5910):

close, but no

OpenStudy (anonymous):

regular 3?

jimthompson5910 (jim_thompson5910):

positive 3, yes

jimthompson5910 (jim_thompson5910):

so k = 3

jimthompson5910 (jim_thompson5910):

y = k*x turns into y = 3*x

jimthompson5910 (jim_thompson5910):

so I'm sure you see the rule now

OpenStudy (anonymous):

either C or D but looks like more like C

jimthompson5910 (jim_thompson5910):

why C?

OpenStudy (anonymous):

wait looks more like D since its 3 times right?

jimthompson5910 (jim_thompson5910):

yep, D is correct

OpenStudy (anonymous):

xDDDDDDD

OpenStudy (anonymous):

haha ty bet help eva btr than my teachers medals and best responses for yu!!!!!